# Spatially resolved H emission in B14-65666: compact starbursts, ionizing efficiency and gas kinematics in an advanced merger at the Epoch of Reionization

### v1

@@ 2.1.2 MIRI imaging
The MIRIM observations of B14-65666 were performed using the F560W filter covering the rest-frame optical continuum at â0.7âm and the H emission line. The MIRIM F560W image has a total on-source integration time of 2675 seconds. The observational setup consists of a 4-point dither pattern, with the dither composed of 2 integrations of 120 groups using the FASTR1 readout mode.
The MIRIM F560W observations are calibrated with version 1.12.0 of the JWST pipeline (Bushouse et al., 2024) and context 1150 of the CRDS. We follow the general procedure to calibrate the data, but we also include more steps to correct for striping artifacts and background gradients (see PĂ©rez-GonzĂĄlez et al. 2024a; Ăstlin et al. 2025 for details). The final MIRIM F560W mosaic has a pixel size of 0.06ââ, after resampling, and a field of view (FoV) of 74âââ113ââ. The Point Spread Function (PSF) Full Width Half Maximum (FWHM) is 0.207ââ (Libralato et al. 2024).
@@ 2.2 NIRCam
The NIRCam (Rieke et al., 2023) data used in this work was presented in a previous study (Sugahara et al. 2025). The observations were taken with filters F115W, F150W, F200W, F277W, F356W and F444W (JWST GO1 program ID 1840; PIs: J. Ălvarez-MĂĄrquez and T. Hashimoto). At redshift 7.15, F115W, F150W and F200W sample the rest-frame ultraviolet spectrum, F277W and F356W cover the Balmer break region, and F444W targets the rest-frame continuum around 5000âĂ and the [OâIII]â4960,5008â emission lines.
NIRCam and MIRIM images present large variation on the spatial resolution, with PSF FWHMs increasing with wavelength from 0.04ââ to 0.21ââ. To obtain consistent multi-wavelength aperture photometry in the different NIRCam and MIRIM images, we perform a PSF homogenization in the NIRCam images to match the spatial resolution of NIRCam images to the coarser MIRIM F560W image. However, we also use the NIRCam images with the native spatial resolution to investigate the structure of B14-65666 in the rest-frame UV (see Sect. 3.4).
The PSF homogenization procedure includes the following steps. We employ the observed effective PSFs (ePSFs) for MIRI (Libralato et al. 2024) and for NIRCam (Nardiello et al. 2022)111https://www.stsci.edu/~jayander/JWST1PASS/LIB/PSFs/STDPSFs/. We compare these ePSFs with point-like sources from our observations, and conclude that they are consistent with each other. The first step is to rotate the NIRCam ePSF to the position angle (PA) of the MIRI image (PA=108Âș) ensuring orientation alignment between the two ePSFs. We rotate the MIRI ePSF to match the PA of the MIRI image. Next, we resample the pixel scale of the MIRI ePSF (0.028ââ, Libralato et al. 2024) to match that of the MIRI imager, and the NIRCam ePSF (0.008ââ for F115W, F150W, F200W and 0.016ââ for F277W, F356W, F444W, Nardiello et al. 2022) to match the pixel scale of the MIRI imager. Using these aligned and resampled ePSFs, we then generate a matching kernel using the python package PHOTUTILS to bridge the differences between the NIRCam and MIRI ePSFs. Finally, to apply the kernel using the python function creatematchingkernel (Bradley et al. 2022), we rescale it to the NIRCam image pixel size (0.03ââ after resampling). Subsequently, we convolve the NIRCam images with the kernel, thus homogenizing the PSF.
@@ 2.3 ALMA ancillary data
The ALMA data used in this work was previously presented in other studies. ALMA Bands 6 and 8 data were taken in Cycles 4 (ID 2016.1.00954.S, PI: A. K. Inoue) and 5 (ID 2017.1.00190.S, PI: A. K. Inoue) to target the [CâII]â158âm and [OâIII]â88âm emission lines (Hashimoto et al. 2019). ALMA Band 7 data was taken during Cycle 7 (ID 2019.1.01491.S, PI: A. K. Inoue; Sugahara et al. 2021). In this article, we use the spectra of [CâII]â158âm and [OâIII]â88âm, and the emission line and continuum (90, 120 and 160âm) fluxes presented in Hashimoto et al. (2019) and Sugahara et al. (2021).
@@ 2.4 MIRI-NIRCam astrometry
We correct the astrometry in the MIRI F560W image by using the centroid of two point-like sources within the FoV of the available GAIA DR3 (Gaia Collaboration et al., 2023) coordinates. The uncertainty in the final positioning is 18 mas. With the corrected MIRIM F560W image, we align the simultaneous MIRIM F770W image with two point-like sources in the common FoV of both images. As the MRS data and the MIRIM F770W image were taken simultaneously, we use the astrometric offset found in the MIRIM F770W image to correct the MRS 1SHORT cube, getting an uncertainty of less than 60âmas (0.3âkpc). Finally, we align the NIRCam images using the astrometrically corrected MIRIM F560W image as for the simultaneous MIRIM F770W image. The NIRCam images have an uncertainty smaller than 35âmas, obtained with the available star and galaxy that have GAIA astrometry (Gaia Collaboration et al., 2023).
@@ 3.1 MIRI and NIRCam images and H line map
Previous HST images showed that B14-65666 is composed of two spatially separated galaxies in the rest-frame UV, a brighter galaxy in the northeast (galaxy E) and a fainter galaxy in the southwest (galaxy W) (Bowler et al. 2017). ALMA data further demonstrated that the [OâIII]â88âm and [CâII]â158âm emission lines can be spatially and spectrally decomposed into two components associated with galaxies E and W (Hashimoto et al. 2018). NIRCam images revealed the complex morphology of the individual galaxy components (Sugahara et al. 2025): galaxy E has a compact core surrounded by diffuse extended optical emission, while galaxy W displays elongated clumpy morphology.
MIRIM imaging and the MRS cover the emission line of H together with the continuum at 0.7âm. MIRI observations show a similar trend as NIRCam images. MRS shows an extended H emission associated with E and W galaxy components, additionally, each of these galaxy components is spectroscopically resolved as already presented in ALMA observations (Hashimoto et al. 2019). We present MIRI and NIRCam images in Fig. 1.
For this study, we define a total elliptical aperture for the whole system (named âgalaxies E+Wâ) with semi-major axis of 0.456ââ and semi-minor axis of 0.343ââ, and position angle (PA) of 85.83Âș. In the case of the galaxy E we define an elliptical aperture with semi-major axis of 0.211ââ and semi-minor axis of 0.201ââ, position angle (PA) of 112.83Âș. For the galaxy W we extract an ellipse with 0.228ââ and 0.177ââ semi-major and semi-minor axis respectively, PA of 112.83Âș. The coordinates of the center are shown in Table 2. The E and W elliptical apertures are compatible with the main emission of these galaxy components in our MIRIM F560W image. âGalaxies E+Wâ elliptical aperture covers the extended H emission from the MRS H line map. The H line map is generated by integrating the H line emission, centered at the H peak of galaxies E+W (5.3510âm) in the velocity range -500âÂĄâvâ[kmâs-1]âÂĄâ300. See Fig. 1.
@@ 3.2 H spectra and line fluxes
We extract the integrated H spectra of galaxies E+W, E and W using the three defined apertures (Fig. 1). We correct the two individual spectra for aperture losses, assuming that the two galaxies are unresolved sources for the MRS angular resolution taking into account the MRS PSF (Argyriou et al. 2023), which has a FWHM of 0.27ââââ0.35ââ at 5.35âm. The aperture correction is 1/0.57 both for galaxies E and W. In the case of the galaxies E+W, we also correct for aperture losses, assuming that the flux is mainly dominated by galaxies E and W, which are unresolved. The aperture correction in this case is 1/0.81. We also correct the spectrum for the contamination of flux from galaxy E in the W aperture by modeling galaxy E at the center of the aforementioned PSF and measuring how much of its flux would appear in the W aperture. We apply the same procedure for galaxy W to estimate its contribution to the E aperture. These corrections are 1/0.02 and 1/0.03, respectively.
@@
We extract twelve 1D spectra using the same apertures as the galaxies E+W, E and W respectively at random positions of the MRS FoV clean of emission. We combine these spectra to generate the 1D median of the local background. The median is subtracted from the H spectra with the aim of removing any systematic residual feature left in the MRS calibration process. The three spectra are presented in Fig. 2.
@@
The H spectrum is modeled by a single Gaussian function, to fit the emission line, and a second-order polynomial, to fit any residual background gradient. Using a Monte Carlo simulation, we estimate the uncertainties for the derived emission line parameters (such as FWHM, flux, and central wavelength). To determine the noise level in the spectrum, the root mean square (rms) is calculated from the continuum adjacent to the emission line. Using this rms value as the 1 standard deviation, we generate 1000 synthetic spectra by adding random Gaussian noise to the original spectrum. Each synthetic spectrum is then refitted to obtain new measurements of the line parameters. The final uncertainty for each parameter is given by the standard deviation of these measurements across all simulations. The Gaussian fit is shown in Fig. 2 and the results are presented in Table 2.
@@
From the H line, the galaxies E+W of B14-65666 is at a redshift of 7.1513ââ0.0007. Galaxy E is redshifted to 7.1529ââ0.0004, however, galaxy W is blueshifted to 7.1481ââ0.0007. This implies that galaxy E and W present a velocity offset between the two components of 175ââ28 kmâs-1, compatible with that observed from [OâIII]â88âm and [CâII]â158âm, 177ââ16 kmâs-1 (Hashimoto et al. 2019).
@@
The H line is spectrally resolved with the MIRI-MRS. It presents an intrinsic FWHM of 407ââ86 kmâs-1 for galaxies E+W and 243ââ41âkmâs-1 and 312ââ44âkmâs-1 respectively for E and W, after deconvolving with the Line Spread Function of the MRS (Labiano et al. 2021). Due to the offset in velocities the galaxies E+W component presents a higher FWHM than galaxy E and W. The observed H fluxes are for E+W, E and W, (23.1ââ3.1)â, (11.9ââ1.3)â, and (8.6ââ1.2)ââerg s-1âcm-2, respectively. All results are presented in Table 2. The H flux for galaxies E+W is in agreement within the uncertainties of the sum of the individual galaxy fluxes. This suggests that the H emission is mainly dominated by the two galaxies excluding any relevant diffuse component.
@@ 3.3 MIRIM and NIRCam photometry
MIRIM F560W photometry for galaxies E+W and for galaxies E and W is derived with the same apertures as for the H spectra (see Fig. 1 and Sect. 3.1). The local background level for each aperture is estimated in an elliptical annulus centered at the same position as the galaxies E+W elliptical aperture and with an inner radius of 0.95ââ and outer radius of 1.95ââ. Additionally, we use 300 apertures on blank regions of the MIRI image with the same size as the galaxy apertures to estimate the flux uncertainties. We also perform the NIRCam photometry using the same aperture as for MIRIM F560W. To do so, we utilize the NIRCam images (F115W, F150W, F200W, F277W, F356W, F444W) that are homogenized to the MIRI F560W PSF. The final observed MIRIM F560W and NIRCam fluxes are presented in Table 2. These NIRCam fluxes are lower by a factor of than the ones obtained from previous result (Sugahara et al. 2025) because of the PSF homogenization from NIRCam F444W to MIRI F560W (see Sect. 2.2) and the different aperture sizes.
Galaxy E is an unresolved source in the NIRCam F150W image with an estimated physical size of 63 pc (see Sect. 3.4 for details). The selected aperture mainly contains emission from galaxy E, thus, we implement aperture correction in the NIRCam and MIRIM photometry using the MIRIM F560W PSF. This aperture correction has a value of 1/0.67. On the contrary, galaxy W is fully resolved in the NIRCam F150W, and for that reason, we do not implement aperture correction to its photometry.
@@ 3.4 Stellar structure and effective radius
We study the stellar structure of the two galaxies, E and W, individually. To achieve this, we utilize the F150W image, as it offers the best combination of signal to noise ratio and spatial resolution. It presents the largest exposure time, 1675âs, arriving at a sensitivity of 7.7ânJy at 5â depth, which is a factor of 2âââ3 better than any other NIRCam band. Also, the PSF FWHM is 0.067ââ, allowing us to resolve structure up to 0.35âkpc at redshift 7.15 (Sugahara et al. 2025). F150W image traces the UV emission, dominated by the young and attenuated stellar population. We work with native-resolution images (before PSF-matching and resampling), and the analysis is done using the code GALFIT (Peng et al. 2002).
@@
Galaxy E consists of a compact core, which appears unresolved, and a non-dominant extended component. To obtain the effective radius of galaxy E, we model the F150W data in GALFIT with a SĂ©rsic profile (with n=0.5, which gives a Gaussian profile) to fit the diffuse emission and a point-like component (using our ePSF) to fit the more compact part of the galaxy (Fig. 3). The surface brightness of galaxy E is compatible with a nuclear, point-like source plus a diffuse emission component at radii larger than 0.05ââ. However, as the core remains unresolved in the F150W image, we extend our analysis to the F115W image, in which the SNR is lower, but the spatial resolution is higher than in the F150W image. The core of the galaxy E also remains unresolved in the F115W image, thus we obtain an upper limit for the effective radius of 63âpc (0.012ââ).
@@
Galaxy W has a clumpy elongated morphology and it is resolved in the F150W filter. It is well modeled by four different clumps. Three of them are unresolved point sources, which we model in GALFIT with point-like components. The other one is a more diffuse clump, associated with the brightest emission, modeled with a point-like and a SĂ©rsic profile (with an index nâ=â0.5) components (Fig. 3). When considering the galaxy W as a single entity and circularizing the profile, the intrinsic effective radius is 348âpc (0.057ââ), after deconvolving with the NIRCam F150W PSF.
@@
The merger nature of B14-65666 is supported by the morphological analysis using the statmorph code (Rodriguez-Gomez et al. 2019). We analyze galaxies E and W separately and find that in both cases, the shape asymmetry, AS â a reliable tracer of merging galaxies (Rinaldi et al. 2024a) â is greater than 0.2, reinforcing the idea that each galaxy exhibits significant asymmetry. Specifically, we find AS of 0.35 and 0.80 for galaxies E and W, respectively. We also test the multimode-intensity-deviation () statistics, which are useful for detecting multi-component systems, such as the double-nucleus of a late-stage merger, highly disordered post-merger remnants, galaxies with bright star-forming clumps in rest-UV emission or an apparently single galaxy with an extended emission component(s) (Freeman et al. 2013). (multimode statistics) quantifies how a galaxyâs light is distributed into distinct regions, and (deviation statistic) measures the offset between the brightest region and the overall centroid. values approaching 1 are likely to indicate a double-nucleus, while values close to zero are interpreted as a single source; values close to zero indicate symmetric and ordered morphologies such as spheroids and disks. In the case of galaxy E, ââ0.0 and ââ0.1. This corroborates the interpretation that galaxy E is a single (compact spheroidal) source. In the case of galaxy W, ââ0.4 and â0.3, indicating that galaxy W is composed of at least two clumps, with one of them dominating the emission.

### v2

@@ 2.1.2 MIRI imaging
The MIRIM observations of B14-65666 were performed using the F560W filter covering the rest-frame optical continuum at â0.7âm and the H emission line. The MIRIM F560W image has a total on-source integration time of 2675 seconds. The observational setup consists of a four-point dither pattern, with the dither composed of two integrations of 120 groups using the FASTR1 readout mode.
The MIRIM F560W observations are calibrated with version 1.12.0 of the JWST pipeline (Bushouse et al., 2024) and context 1150 of the CRDS. We followed the general procedure to calibrate the data, but we also included more steps to correct for striping artifacts and background gradients (see PĂ©rez-GonzĂĄlez et al. 2024a; Ăstlin et al. 2025 for details). The final MIRIM F560W mosaic has a pixel size of 0.06ââ, after resampling, and a field of view (FoV) of 74âââ113ââ. The point spread function (PSF) full width at half maximum (FWHM) is 0.207ââ (Libralato et al. 2024).
@@ 2.2 NIRCam
The NIRCam (Rieke et al., 2023) data used in this work were presented in a previous study (Sugahara et al. 2025). The observations were taken with filters F115W, F150W, F200W, F277W, F356W, and F444W (JWST GO1 program ID 1840; PIs: J. Ălvarez-MĂĄrquez and T. Hashimoto). At redshift 7.15, F115W, F150W, and F200W sample the rest-frame UV spectrum, F277W and F356W cover the Balmer break region, and F444W targets the rest-frame continuum around 5000âĂ and the [OâIII]â4960,5008â emission lines.
NIRCam and MIRIM images present large variations on the spatial resolution, with PSF FWHMs increasing with wavelength from 0.04ââ to 0.21ââ. To obtain consistent multiwavelength aperture photometry in the different NIRCam and MIRIM images, we performed a PSF homogenization in the NIRCam images to match the spatial resolution of NIRCam images to the coarser MIRIM F560W image. However, we also used the NIRCam images with the native spatial resolution to investigate the structure of B14-65666 in the rest-frame UV (see Sect. 3.4).
The PSF homogenization procedure includes the following steps. We employed the observed effective PSFs (ePSFs) for MIRI (Libralato et al. 2024) and for NIRCam (Nardiello et al. 2022)111https://www.stsci.edu/~jayander/JWST1PASS/LIB/PSFs/STDPSFs/.. We compared these ePSFs with point-like sources from our observations and conclude that they are consistent with each other. The first step was to rotate the NIRCam ePSF to the position angle (PA) of the MIRI image (PA=108Âș), ensuring orientation alignment between the two ePSFs. We rotate the MIRI ePSF to match the PA of the MIRI image. Next, we resampled the pixel scale of the MIRI ePSF (0.028ââ, Libralato et al. 2024) to match that of the MIRI imager and the NIRCam ePSF (0.008ââ for F115W, F150W, and F200W; and 0.016ââ for F277W, F356W, and F444W; Nardiello et al. 2022) to match the pixel scale of the MIRI imager. Using these aligned and resampled ePSFs, we then generated a matching kernel using the python package PHOTUTILS to bridge the differences between the NIRCam and MIRI ePSFs. Finally, to apply the kernel using the python function creatematchingkernel (Bradley et al. 2022), we rescaled it to the NIRCam image pixel size (0.03ââ after resampling). Subsequently, we convolved the NIRCam images with the kernel, thus homogenizing the PSF.
@@ 2.3 ALMA ancillary data
The ALMA data used in this work were previously presented in other studies. ALMA Band-6 and Band-8 data were taken in Cycles 4 (ID 2016.1.00954.S, PI: A. K. Inoue) and 5 (ID 2017.1.00190.S, PI: A. K. Inoue) to target the [CâII]â158âm and [OâIII]â88âm emission lines (Hashimoto et al. 2019). ALMA Band-7 data were taken during Cycle 7 (ID 2019.1.01491.S, PI: A. K. Inoue; Sugahara et al. 2021). In this article, we used the spectra of [CâII]â158âm and [OâIII]â88âm and the emission-line and continuum (90, 120, and 160âm) fluxes presented in Hashimoto et al. (2019) and Sugahara et al. (2021).
@@ 2.4 MIRI-NIRCam astrometry
We corrected the astrometry in the MIRI F560W image by using the centroid of two point-like sources within the FoV of the available GAIA DR3 (Gaia Collaboration et al., 2023) coordinates. The uncertainty in the final positioning is 18 mas. With the corrected MIRIM F560W image, we aligned the simultaneous MIRIM F770W image with two point-like sources in the common FoV of both images. As the MRS data and the MIRIM F770W image were taken simultaneously, we used the astrometric offset found in the MIRIM F770W image to correct the MRS 1SHORT cube, obtaining an uncertainty of less than 60âmas (0.3âkpc). Finally, we aligned the NIRCam images using the astrometrically corrected MIRIM F560W image as we did for the simultaneous MIRIM F770W image. The NIRCam images have an uncertainty smaller than 35âmas; this was obtained with the available star and galaxy that have Gaia astrometry (Gaia Collaboration et al., 2023).
@@ 3.1 MIRI and NIRCam images and H line map
Previous HST images showed that B14-65666 is composed of two spatially separated galaxies in the rest-frame UV, a brighter galaxy in the northeast (galaxy E), and a fainter galaxy in the southwest (galaxy W) (Bowler et al. 2017). ALMA data further demonstrated that the [OâIII]â88âm and [CâII]â158âm emission lines can be spatially and spectrally decomposed into two components associated with galaxies E and W (Hashimoto et al. 2018). NIRCam images revealed the complex morphology of the individual galaxy components (Sugahara et al. 2025): galaxy E has a compact core surrounded by diffuse extended optical emission, while galaxy W displays elongated clumpy morphology.
The MIRIM and MRS data cover the emission line of H together with the continuum at 0.7âm. MIRI observations show a similar trend to NIRCam images. MRS shows an extended H emission associated with E and W galaxy components; additionally, each of these galaxy components is spectroscopically resolved as they were already presented in ALMA observations (Hashimoto et al. 2019). We present MIRI and NIRCam images in Fig. 1.
For this study, we defined a total elliptical aperture for the whole system (named galaxies E+W) with a semi-major axis of 0.456ââ, semi-minor axis of 0.343ââ, and position angle (PA) of 85.83Âș. In the case of the galaxy E we define an elliptical aperture with semi-major axis of 0.211ââ and semi-minor axis of 0.201ââ, position angle (PA) of 112.83Âș. For the galaxy W, we extracted an ellipse with a 0.228ââ and 0.177ââ semi-major and semi-minor axis, respectively, and a PA of 112.83Âș. The coordinates of the center are shown in Table 2. The E and W elliptical apertures are compatible with the main emission of these galaxy components in our MIRIM F560W image. The elliptical aperture of galaxies E+W covers the extended H emission from the MRS H line map. The H line map is generated by integrating the H line emission, centered at the H peak of galaxies E+W (5.3510âm) in the velocity range -500âÂĄâvâ[kmâs-1]âÂĄâ300 (see Fig. 1).
@@ 3.2 H spectra and line fluxes
We extracted the integrated H spectra of galaxies E+W, E, and W using the three defined apertures (Fig. 1). We corrected the two individual spectra for aperture losses, assuming that the two galaxies are unresolved sources for the MRS angular resolution taking into account the MRS PSF (Argyriou et al. 2023), which has a FWHM of 0.27ââââ0.35ââ at 5.35âm. The aperture correction is 1/0.57 for the galaxies E and W. In the case of the galaxies E+W, we also corrected for aperture losses, assuming that the flux is mainly dominated by galaxies E and W, which are unresolved. The aperture correction in this case is 1/0.81. We also corrected the spectrum for the contamination of flux from galaxy E in the W aperture by modeling galaxy E at the center of the aforementioned PSF and measuring how much of its flux would appear in the W aperture. We applied the same procedure for galaxy W to estimate its contribution to the E aperture. These corrections are 1/0.02 and 1/0.03, respectively.
@@
We extracted twelve 1D spectra using the same apertures as the E+W, E, and W galaxies, respectively, at random positions of the MRS FoV clean of emission. We combined these spectra to generate the 1D median of the local background. The median is subtracted from the H spectra with the aim of removing any systematic residual feature left in the MRS calibration process. The three spectra are presented in Fig. 2.
@@
The H spectrum is modeled by a single Gaussian function, to fit the emission line, and a second-order polynomial, to fit any residual background gradient. Using a Monte Carlo simulation, we estimated the uncertainties for the derived emission line parameters (such as FWHM, flux, and central wavelength). To determine the noise level in the spectrum, the root mean square (rms) was calculated from the continuum adjacent to the emission line. Using this rms value as the 1 standard deviation, we generated 1000 synthetic spectra by adding random Gaussian noise to the original spectrum. Each synthetic spectrum was then refit to obtain new measurements of the line parameters. The final uncertainty for each parameter is given by the standard deviation of these measurements across all simulations. The Gaussian fit is shown in Fig. 2, and the results are presented in Table 2.
@@
From the H line, the E+W galaxies of B14-65666 are at a redshift of 7.1513ââ0.0007. Galaxy E is redshifted to 7.1529ââ0.0004; however, galaxy W is blueshifted to 7.1481ââ0.0007. This implies that galaxy E and W present a velocity offset between the two components of 175ââ28 kmâs-1, compatible with that observed from [OâIII]â88âm and [CâII]â158âm, 177ââ16 kmâs-1 (Hashimoto et al. 2019).
@@
The H line is spectrally resolved with the MIRI-MRS. It presents an intrinsic FWHM of 407ââ86 kmâs-1 for galaxies E+W and 243ââ41âkmâs-1 and 312ââ44âkmâs-1, respectively, for E and W, after deconvolving with the line spread function of the MRS (Labiano et al. 2021). Due to the offset in velocities the component of the E+W galaxies presents a higher FWHM than that of galaxies E and W. The observed H fluxes are for E+W, E, and W: (23.1ââ3.1)â, (11.9ââ1.3)â, and (8.6ââ1.2)ââerg s-1âcm-2, respectively. All results are presented in Table 2. The H flux for galaxies E+W is in agreement within the uncertainties of the sum of the individual galaxy fluxes. This suggests that the H emission is mainly dominated by the two galaxies excluding any relevant diffuse component.
@@ 3.3 MIRIM and NIRCam photometry
MIRIM F560W photometry for galaxies E+W and for galaxies E and W is derived with the same apertures as for the H spectra (see Fig. 1 and Sect. 3.1). The local background level for each aperture is estimated in an elliptical annulus centered at the same position as the E+W galaxiesâ elliptical aperture and with an inner radius of 0.95ââ and outer radius of 1.95ââ. Additionally, we used 300 apertures on blank regions of the MIRI image with the same size as the galaxy apertures to estimate the flux uncertainties. We also performed the NIRCam photometry using the same aperture as for MIRIM F560W. To do so, we utilized the NIRCam images (F115W, F150W, F200W, F277W, F356W, F444W) that are homogenized to the MIRI F560W PSF. The final observed MIRIM F560W and NIRCam fluxes are presented in Table 2. These NIRCam fluxes are lower by a factor of than the ones obtained from previous results (Sugahara et al. 2025) because of the PSF homogenization from NIRCam F444W to MIRI F560W (see Sect. 2.2) and the different aperture sizes.
Galaxy E is an unresolved source in the NIRCam F150W image with an estimated physical size of 63 pc (see Sect. 3.4 for details). The selected aperture mainly contains emission from galaxy E; thus, we implemented aperture correction in the NIRCam and MIRIM photometry using the MIRIM F560W PSF. This aperture correction has a value of 1/0.67. On the contrary, galaxy W is fully resolved in the NIRCam F150W, and for that reason we did not implement aperture correction to its photometry.
@@ 3.4 Stellar structure and effective radius
We study the stellar structure of the two galaxies, E and W, individually. To achieve this, we utilized the F150W image, as it offers the best combination of signal-to-noise ratio (S/N) and spatial resolution. It presents the largest exposure time, 1675âs, arriving at a sensitivity of 7.7ânJy at 5â depth, which is a factor of 2âââ3 better than any other NIRCam band. Also, the PSF FWHM is 0.067ââ, allowing us to resolve structure up to 0.35âkpc at redshift 7.15 (Sugahara et al. 2025). The F150W image traces the UV emission, which is dominated by the young and attenuated stellar population. We worked with native-resolution images (before PSF-matching and resampling), and the analysis was done using the GALFIT code (Peng et al. 2002).
@@
Galaxy E consists of a compact core, which appears unresolved, and a nondominant extended component. To obtain the effective radius of galaxy E, we modeled the F150W data in GALFIT with a SĂ©rsic profile (with n=0.5, which gives a Gaussian profile) to fit the diffuse emission and a point-like component (using our ePSF) to fit the more compact part of the galaxy (Fig. 3). The surface brightness of galaxy E is compatible with a nuclear, point-like source plus a diffuse emission component at radii larger than 0.05ââ. However, as the core remains unresolved in the F150W image, we extend our analysis to the F115W image, in which the S/N is lower, but the spatial resolution is higher than in the F150W image. The core of galaxy E also remains unresolved in the F115W image; thus, we obtained an upper limit for the effective radius of 63âpc (0.012ââ).
@@
Galaxy W has a clumpy elongated morphology and is resolved in the F150W filter. It is well modeled by four different clumps. Three of them are unresolved point sources, which we modeled in GALFIT with point-like components. The other one is a more diffuse clump, associated with the brightest emission, modeled with point-like and SĂ©rsic-profile (with an index nâ=â0.5) components (Fig. 3). When considering galaxy W as a single entity and circularizing the profile, the intrinsic effective radius is 348âpc (0.057ââ) after deconvolving with the NIRCam F150W PSF.
@@
The merger nature of B14-65666 is supported by the morphological analysis using the statmorph code (Rodriguez-Gomez et al. 2019). We analyzed galaxies E and W separately and find that in both cases the shape asymmetry, AS â a reliable tracer of merging galaxies (Rinaldi et al. 2024a) â is greater than 0.2, reinforcing the idea that each galaxy exhibits significant asymmetry. Specifically, we find AS of 0.35 and 0.80 for galaxies E and W, respectively. We also tested the multimode-intensity-deviation (MID) statistics, which are useful for detecting multicomponent systems, such as the double-nucleus of a late-stage merger, highly disordered post-merger remnants, galaxies with bright star-forming clumps in rest-UV emission, or an apparently single galaxy with an extended emission component(s) (Freeman et al. 2013). M (multimode statistics) quantifies how a galaxyâs light is distributed into distinct regions, and D (deviation statistic) measures the offset between the brightest region and the overall centroid. M values approaching 1 are likely to indicate a double nucleus, while values close to zero are interpreted as indicating a single source; D values close to zero indicate symmetric and ordered morphologies such as spheroids and disks. In the case of galaxy E, Mââ0.0 and Dââ0.1. This corroborates the interpretation that galaxy E is a single (compact spheroidal) source. In the case of galaxy W, Mââ0.4 and D â0.3, indicating that galaxy W is composed of at least two clumps, with one of them dominating the emission.